OpenStudy (anonymous):

Help! If f(x)= 6−5x, the what is f(4)?

OpenStudy (anonymous):

replace your x by 4

OpenStudy (alexandervonhumboldt2):

f(x)=6-5*4

OpenStudy (anonymous):

plug in 4 for x same as last time

OpenStudy (anonymous):

-14?

OpenStudy (anonymous):

6-20= -14 so yes

OpenStudy (anonymous):

How would I set up this one? f(x)= −2x + 15 and f(x)= −11, find x.

OpenStudy (sleepyjess):

If f(x) = -2x + 15 and f(x) = -11, then these are both equal to each other, so, -11 = -2x + 15

OpenStudy (anonymous):

How I would I solve that?

OpenStudy (sleepyjess):

First, how do we get + 15 to the other side?